OpenStudy (anonymous):

what is the inverse of the function y=2x-3?

OpenStudy (anonymous):

i change y with x and x with y, so x=yx+3, and y=2x+3, set to 0?

OpenStudy (mayankdevnani):

y = 2x - 3 interchange x and y x = 2y - 3 now we have to make y the subject of the formula x + 3 = 2y (x + 3) / 2 = y y = (x + 3) / 2 y = 1/2x + 3/2

OpenStudy (mayankdevnani):

ok @Life

OpenStudy (mayankdevnani):

right @hartnn

hartnn (hartnn):

yes, correct.

OpenStudy (mayankdevnani):

thnx